### v2

@@ 5.4 Monte Carlo simulation results
Figure 17: This figure shows the results of our MC analysis. All the plots shown are combined results of 200 simulated samples. The first column of histograms shows the distribution of properties for the 14 sample cases. The second and third columns of histograms present the distribution of properties for 28 and 42 sample cases, respectively. The first row shows the correlation distribution, and the red line shows the mean. The second row shows the RMSE distribution. The third and fourth rows show the bias and NMAD distributions, respectively.
@@
5.4.1 Future of Multi-wavelength Redshift Estimates via Machine Learning
@@
The following subsection draws upon the results from Sec. 5.4 and Sec. 5.3 for estimates on the number of GRB redshifts that can be predicted using ML analysis for future X-ray and optical observations of afterglows. These estimates are based on the estimates presented in (Dainotti et al., 2022b), which in turn rely on the existing X-ray and optical samples determined by Dainotti et al. (2024c) and Dainotti et al. (2024b), provided up to May 2025.
@@
Table 1: The table details the number of GRBs this study will enable to have redshift predictions in the future using Swift observations.
@@
For the X-ray afterglow, we currently have 255 GRBs possessing plateaus and measured $z_{obs}$, observed from 03/2005 to 12/2023. We have 299 events with X-ray plateaus but without redshifts, observed from 01/2005 to 02/2024. Both sets of GRBs are taken from the Swift/XRT catalogs. Based on these data, $\sim$13.6 GRBs with X-ray plateaus and $z_{obs}$, and $\sim$15.7 GRBs with X-ray plateaus and unknown redshifts are expected annually. For GRBs with $z_{obs}$ and X-ray plateau, 24 additional GRBs are expected from 12/2023 until 09/2025. For GRBs with X-ray plateaus and without $z_{obs}$, 25 additional GRBs are also expected from 02/2024 until 09/2025. This sample with unknown redshift will be the generalization set. So, the number of GRBs with $z_{pred}$ from ML analysis will be $>$324.
@@
For the optical afterglow, we have gathered 536 LCs through the 416 ground-based Telescopes, the HST, the Swift/UVOT satellites, the GCN, and private communications. These 536 GRBs, observed from 05/1997 to 10/2023, have $z_{obs}$ and optical afterglow. Out of these 536 GRBs, 180 (34%) GRBs have optical plateau observations. As of this writing, according to the Greiner webpage [^3] https://www.mpe.mpg.de/ jcg/grbgen.html, the total number of GRBs with optical afterglow is 932, and 622 have secure redshifts. Thus, 310 GRBs (932 - 622) do not have redshift but have an optical afterglow. Assuming the same fraction of plateau observations in the optical afterglow, our current generalization set would be 34% of 310, which is 104 GRBs. If we extrapolate to 09/2025, we expect to have 112 GRBs with an optical plateau and no measured redshifts. Similarly, we expect to have 193 GRBs by 09/2025, which have an optical plateau and $z_{obs}$. Thus, we expect $>$112 GRBs with $z_{pred}$ from ML analysis. Thus, the new $z_{pred}$ will be $>$436 (324+112), where 324 and 112 are GRBs with X-ray and optical plateaus, respectively. With this increase of 92% in the redshift sample, we will enhance the accuracy of the LF and density rate evolution estimates, thus helping to address the discrepancy between the GRBFR and SFR evolution. The above numbers are also reflected in the Table 1. This analysis can be extended for future reference, and it is our basis to compute how many GRBs we will have approximately in the next one, two, and three years from now.
@@
According to Table.9 of Dainotti et al. (2022), to reach the precision of Conley et al. (2011), the lowest number of X-ray platinum sample GRBs with redshift needed, with half error bars on plateau parameters, is 354. To achieve this number, we assume that by using an updated lightcurve reconstruction (LCR) technique, an improvement of the methodology from Dainotti et al. (2023), which is in preparation, we will have the same number of GRBs with half error bars (denoted by n=2) as with full error bars (denoted by n=1). Currently, we have 105 GRBs that belong to the platinum sample with $z_{obs}$. Assuming the same fraction of platinum GRBs exists in the generalization set, we should have another 125 GRBs from the generalization set used in this work. Combining these two, we estimate that we currently have 230 platinum sample GRBs with pseudo-redshifts. Further, considering the observations we will obtain from SVOM (Godet et al. (2014) and SVOM webpage [^4] https://www.eoportal.org/satellite-missions/svom-eclairs-gamma-ray-imager) and the application of ML and LCR, we can expect to reach the precision of Conley et al. (2011) by 2029 (see PLATtrim20 column of Table.2). Thus, this ML effort reduces the number of years required to reach the precision of Conley et al. (2011) by 26 years. Similarly, to reach the precision of Betoule et al. (2014), we need the observational capabilities of THESEUS. According to our estimates in Table.2, again, using LCR and ML to boost the numbers and GRBs with half error, we should be able to reach the precision of Betoule et al. (2014) by 2041 at least.
@@
Similarly, let us consider the optical estimates from Table 9 of Dainotti et al. (2022). We should be able to reach the precision of Conley et al. (2011) with the current data we have if we assume that we will have an equal number of half error GRBs as we have full error GRBs, achieved by using updated LCR methodology. If we, however, trim the sample in the closest 25 to the 3D fundamental plane (OPTtrim25 column of Table.3), we would have reached the precision of Conley et al. (2011) even in 2005, 6 years earlier than the one reached by the SNe Ia. Note that the numbers mentioned along the fourth row of Table.3 are estimated based on GRBs observed from January 2005. This is done to avoid negative numbers, as some of the estimates are reached well before 2024. The corresponding year estimates are obtained based on the rate of observation using Swift and other ground-based optical observatories. Furthermore, if we take into account the rate of observation of SVOM and the fraction of optical GRBs with the plateau we currently have, then we can reach the precision of Conley et al. (2011) by 2028 (only 4 years from now), with the current errorbars. However, to reach the precision of Betoule et al. (2014), we need to consider the observation of THESEUS as well. Considering THESEUS’s rate of observation to be 700 GRBs per year (Dainotti et al., 2022; Amati et al., 2018a), we again consider the same fraction of optical plateau GRBs as we currently have. With this, we can reach the precision of Betoule et al. (2014) by 2039, 7 years after the launch of THESEUS (the current expected launch is 2032, if the mission is approved). However, to reach the same precision as (Scolnic et al., 2018), we would need to wait at least until 2041 and use both ML and LCR and half error bars.
@@
However, we can say that these estimates are rather conservative since new optical telescopes and facilities will be operating in the future, and we can rely on the efforts of some of us to continue building the largest optical GRB catalog to date (Dainotti et al., 2024a). Indeed, one of the outcomes of gathering multiple data points from several sources has the advantage of allowing an increase of data points, in general, in the LCs and, in particular, in the plateau region when it occurs. Thus, we assume that once the optical GRB sample is updated, our estimates will be larger in terms of the total number of optical plateaus, and thus, the number of years needed to reach SNe Ia precision will be less. In addition, regarding the X-ray sample, some of us are currently working on choosing a subsample, not driven by phenomenological reasons but driven by the magnetar model (Zhang & Mészáros, 2001; Rowlinson et al., 2014b; Yi et al., 2017). In such a case, we can have a smaller scatter of the plane as shown for the 2D Dainotti relation between the X-ray luminosity at the end of the plateau and its rest-frame duration Dainotti et al. (2008, 2011); Dainotti213; ainotti2017; Wang et al. (2022), and we can reach again the precision of the SNe Ia sooner, using the X-ray sample.
Table 2: This table presents the number of platinum sample Xray GRBs with redshift needed to reach the precision of $\Omega_{M}$ as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia). This is an extension of the table provided in Dainotti et al. (2022). n=1 indicates that the error bars on the GRB plateau parameters are full-width, while n=2 indicates that the error bars are halved. LCR indicates that the estimates for the number of GRBs have been boosted by using LCR, while ML indicates that the estimates have been boosted by using machine learning.
@@
Table 3: This table presents the number of optical GRBs with plateau and redshift needed to reach the precision of $\Omega_{M}$ as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia). This is an extension of the table provided in Dainotti et al. (2022). n=1 indicates that the error bars on the GRB plateau parameters are full-width, while n=2 indicates that the error bars are halved. LCR indicates that the estimates for the number of GRBs have been boosted by using LCR, while ML indicates that the estimates have been boosted by using machine learning.
@@ 5.5 Applicability to future missions
It is important to note that all the methodologies spelled out here can be applied not only to GRBs observed by Swift, but also to any GRB for which we can observe the plateau emission. Thus, we envision this method can be applied to GRBs observed by Space-based multi-band astronomical Variable Objects Monitor (SVOM,Wei et al. (2016)), the Transient High-Energy Sky and Early Universe Surveyor (THESEUS) mission (Amati et al., 2018b), and the Einstein Probe mission(Yuan et al., 2015). Surely, with the new data from other satellites, one needs to be cautious about how the variables’ parameter spaces will change compared to the space of the Swift variables and how much this will impact the prediction. Properly leveraging the power of these new satellites will require a re-training of the ML models. However, the base methodology described here and in our previous works (Dainotti et al., 2024c, b) should still produce reliable results.
@@ 5.6 Redshift inference with the web app
Our web app has various functionalities, but its main focus is inferring the redshift of GRBs without $z_{obs}$. Users have the flexibility to upload a file containing the predictors discussed in this work (see Sec. 4.2.1) or manually enter the parameters of a single GRB for redshift estimation. This interface is presented in Fig. 18. This module of the web app also allows the user to obtain MICE imputations if their data contains missing values. To do so, they select the ‘Apply MICE’ button. After providing the data, the web app initiates the back-end ML codes, the results and performance of which have been presented in this work. Upon successful execution, the user is presented with individual redshift values for each GRB in a tabular manner, which can also be downloaded as a text file. This table displays the GRB IDs along with $z_{pred}$ and its associated uncertainties. Furthermore, if a user already has their custom model, trained from modules such as Model Comparision and SuperLearner, then they can use that model for predictions. For this, they need to select the ‘Do you have modeled data?’ option. However, the users can use the model trained on our data without necessarily providing their own training set. Finally, comprehensive documentation is available on the web app in the ‘Documentation’ module to help users navigate.
Figure 18: The figure shows the homepage of the app, which is the Redshift Estimator. It also shows the panel where the user needs to upload the data file, the option for applying MICE and using custom models.
@@ 6 Discussion
The findings from Dainotti et al. (2024b) demonstrate the promise of using machine learning for redshift estimation, particularly in overcoming the limited number of GRBs with precise redshift measurements. In this work, we have expanded both the methodology and the sample sizes for the training and generalization datasets. Providing redshift predictions for the generalization set will help future population studies address the GRBFR and SFR discrepancy and other relevant issues. In the future, the $z_{pred}$ values obtained here can also be utilized to explore the characteristics of the central engine in GRB events, focusing on GRBFR, LF, and the structure of LCs. We can also use GRB LCs obtained from reconstruction, which some of us did in Dainotti et al. (2023). We demonstrated the effectiveness of this reconstruction technique in reducing the errors on plateau parameters by an average of 30%-40%.
Another point we would like to highlight here is that even though we have the largest sample of LGRBs with plateau and redshift, this is still at least an order of magnitude less than the data required to reliably train more complex ML models, such as neural networks. Indeed, without including the plateau, Swift XRT observed 1420 GRBs up to December 2024, which is too small to train complex ML models properly. In some of our previous works, using Active Galactic Nuclei data (Dainotti et al., 2021a; Narendra et al., 2022; Gibson et al., 2022), we have shown that SuperLearner picks different, more complex models (e.g., XGBoost, EARTH) for the larger sample size. Similarly, we tested multiple ML models to find the best set of models that can give accurate redshift predictions with the current data set. The three chosen models performed the best during our analysis. Given the size of our data set, parametric and semi-parametric functions like GLM and GAM perform better. This fact is further highlighted in the Appendix in Fig. 20, where the weight assigned by SuperLearner to the model EARTH is lower than the weights assigned to the GAM, GLM, and Random forest models. This is in contrast with the results presented in Narendra et al. (2022) where Earth was one of the best models. Furthermore, to test whether the performance of these models is maintained as the dataset increases, we have also performed a Monte Carlo simulation (see Sect. 5.4). From those simulations we can say that our models will be reliable for up to 3 years, after which we will require to retrain them with different models and training data. Thus, based on our analysis, these chosen models are best suited for the data at hand.
6.1 LGRBs vs SGRBs

### v3

@@ 5.4 Monte Carlo simulation results
Figure 17: Results of our MC analysis. All the plots shown are combined results of 200 simulated samples. The first column of histograms shows the distribution of properties for the 14 sample cases. The second and third columns of histograms present the distribution of properties for 28 and 42 sample cases, respectively. The first row shows the correlation distribution, and the red line shows the mean. The second row shows the RMSE distribution. The third and fourth rows show the bias and NMAD distributions, respectively.
@@
5.5 Future of multi-wavelength redshift estimates via machine learning
@@
The following subsection draws upon the results from Sec. 5.4 and Sec. 5.3 for estimates on the number of GRB redshifts that can be predicted using ML analysis for future X-ray and optical observations of afterglows. These estimates are based on the estimates presented in (Dainotti et al., 2023a), which in turn rely on the existing X-ray and optical samples determined by Dainotti et al. (2024c) and Dainotti et al. (2024b), provided up to May 2025.
Table 1: Present and future numbers of GRB observations.
GRB subsets Number Current data gathered from-to Number of yrs Rate/yr Data gathering until September 2025 GRBs by September 2025
Optical Afterglow with z 536 05/1997 - 10/2023 27 19.7 10/2023-09/2025 574
Optical plateau with z 180 05/1997-05/2021 24 7.50 05/2021-09/2025 193
Optical Afterglow without z 310 05/1997- 10/2023 26 11.92 10/2023-09/2025 333
Optical plateau without z 104 05/1997- 10/2023 26 4.00 10/2023-09/2025 112
X-ray plateaus without z 299 01/2005 - 02/2024 19.1 15.7 02/2024-09/2025 324
X-ray plateaus with z 255 03/2005 - 12/2023 18.7 13.6 12/2023-09/2025 279
Known redshifts in X-rays & Optical by May 2025 279 (X-ray) + 193(OA) From May 1997 to September 2025 472
Predicted redshift via ML in X-rays and Optical by May 2025 324 (X-ray) + 112 (OA) 436
% increase from X-rays and Optical plateaus with z with ML 436/472 92% more redshifts will be estimated
@@
 [^3] Number of GRBs this study will enable to have redshift predictions in the future using Swift observations.
@@
For the X-ray afterglow, we currently have 255 GRBs possessing plateaus and measured $z_{obs}$, observed from 03/2005 to 12/2023. We have 299 events with X-ray plateaus but without redshifts, observed from 01/2005 to 02/2024. Both sets of GRBs are taken from the Swift/XRT catalogs. Based on these data, $\sim$13.6 GRBs with X-ray plateaus and $z_{obs}$, and $\sim$15.7 GRBs with X-ray plateaus and unknown redshifts are expected annually. For GRBs with $z_{obs}$ and X-ray plateau, 24 additional GRBs are expected from 12/2023 until 09/2025. For GRBs with X-ray plateaus and without $z_{obs}$, 25 additional GRBs are also expected from 02/2024 until 09/2025. This sample with unknown redshift will be the generalization set. So, the number of GRBs with $z_{pred}$ from ML analysis will be greater than 324.
@@
For the optical afterglow, we have gathered 536 LCs through the 416 ground-based Telescopes, the HST, the Swift/UVOT satellites, the GCN, and private communications. These 536 GRBs, observed from 05/1997 to 10/2023, have $z_{obs}$ and optical afterglow. Out of these 536 GRBs, 180 (34%) GRBs have optical plateau observations. As of this writing, according to the Greiner webpage [^4] https://www.mpe.mpg.de/ jcg/grbgen.html, the total number of GRBs with optical afterglow is 932, and 622 have secure redshifts. Thus, 310 GRBs (932 - 622) do not have redshift but have an optical afterglow. Assuming the same fraction of plateau observations in the optical afterglow, our current generalization set would be 34% of 310, which is 104 GRBs. If we extrapolate to 09/2025, we expect to have 112 GRBs with an optical plateau and no measured redshifts. Similarly, we expect to have 193 GRBs by 09/2025, which have an optical plateau and $z_{obs}$. Thus, we expect more than 112 GRBs with $z_{pred}$ from ML analysis. Thus, the new $z_{pred}$ will be more than 436 (324+112), where 324 and 112 are GRBs with X-ray and optical plateaus, respectively. With this increase of 92% in the redshift sample, we will enhance the accuracy of the LF and density rate evolution estimates, and thus it will help address the discrepancy between the GRBFR and SFR evolution. The above numbers are also reflected in the Table 3. This analysis can be extended for future reference, and it serves as our foundation to compute how many GRBs we will approximately have in the next one, two, and three years from now.
@@
According to Table.9 of Dainotti et al. (2022), to reach the precision of Conley et al. (2011), the lowest number of X-ray platinum sample GRBs with redshift needed, with half error bars on plateau parameters, is 354. To achieve this number, we assume that by using an updated LC reconstruction (LCR) technique, an improvement of the methodology from Dainotti et al. (2023b), which is in preparation, we will have the same number of GRBs with half error bars (denoted by n=2) as with full error bars (denoted by n=1). Currently, we have 105 GRBs that belong to the platinum sample with $z_{obs}$. Assuming the same fraction of platinum GRBs exists in the generalization set, we should have another 125 GRBs from the generalization set used in this work. Combining these two, we estimate that we currently have 230 platinum sample GRBs with pseudo-redshifts. Further, considering the observations we will obtain from SVOM (Godet et al. (2014) and SVOM webpage [^5] https://www.eoportal.org/satellite-missions/svom-eclairs-gamma-ray-imager) and the application of ML and LCR, we can expect to reach the precision of Conley et al. (2011) by 2029 (see PLATtrim20 column of Table.6). Thus, this ML effort reduces the number of years required to reach the precision of Conley et al. (2011) by 26 years. Similarly, to reach the precision of Betoule et al. (2014), we need the observational capabilities of THESEUS. According to our estimates in Table.6, again, using LCR and ML to boost the numbers and GRBs with half error, we should be able to reach the precision of Betoule et al. (2014) by 2041 at least.
@@
Similarly, let us consider the optical estimates from Table 9 of Dainotti et al. (2022). We should be able to reach the precision of Conley et al. (2011) with the current data we have if we assume that we will have an equal number of half error GRBs as we have full error GRBs, achieved by using updated LCR methodology. If we, however, trim the sample in the closest 25 to the 3D fundamental plane (OPTtrim25 column of Table.7), we would have reached the precision of Conley et al. (2011) even in 2005, 6 years earlier than the one reached by the SNe Ia. We note that the numbers mentioned along the fourth row of Table.7 are estimated based on GRBs observed from January 2005. This is done to avoid negative numbers, as some of the estimates are reached well before 2024. The corresponding year estimates are obtained based on the rate of observation using Swift and other ground-based optical observatories. Furthermore, if we take into account the rate of observation of SVOM and the fraction of optical GRBs with the plateau we currently have, then we can reach the precision of Conley et al. (2011) by 2028 (only 4 years from now), with the current error bars. However, to reach the precision of Betoule et al. (2014), we need to consider the observation of THESEUS as well. Considering THESEUS’s rate of observation to be 700 GRBs per year (Dainotti et al., 2022; Amati et al., 2018), we again consider the same fraction of optical plateau GRBs as we currently have. With this, we can reach the precision of Betoule et al. (2014) by 2039, 7 years after the launch of THESEUS (the current expected launch is 2032, if the mission is approved). However, to reach the same precision as (Scolnic et al., 2018), we would need to wait at least until 2041 and use both ML and LCR and half error bars.
@@
However, we can say that these estimates are rather conservative since new optical telescopes and facilities will be operating in the future, and we can rely on the efforts of some of us to continue building the largest optical GRB catalog to date (Dainotti et al., 2024a). Indeed, one of the outcomes of gathering multiple data points from several sources has the advantage of allowing an increase of data points, in general, in the LCs and, in particular, in the plateau region when it occurs. Thus, we assumed that once the optical GRB sample is updated, our estimates will be larger in terms of the total number of optical plateaus, and therefore the number of years needed to reach SNe Ia precision will be less. In addition, regarding the X-ray sample, some of us are currently working on choosing a subsample, not driven by phenomenological reasons but driven by the magnetar model (Zhang & Mészáros, 2001; Rowlinson et al., 2014b; Yi et al., 2017). In such a case, we can have a smaller scatter of the plane as shown for the 2D Dainotti relation between the X-ray luminosity at the end of the plateau and its rest-frame duration Dainotti et al. (2008, 2011, 2013, 2017); Wang et al. (2022), and we can reach again the precision of the SNe Ia sooner, using the X-ray sample.
Table 2: Forecast on the future precision of X-ray GRB cosmology.
@@
 [^6] Number of platinum sample X-ray GRBs with redshift needed to reach the precision of $\Omega_{M}$ as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia). This is an extension of the table provided in Dainotti et al. (2022). The term n=1 indicates that the error bars on the GRB plateau parameters are full width, while n=2 indicates that the error bars are halved. The ”LCR” indicates that the estimates for the number of GRBs have been boosted by using LCR, while ”ML” indicates that the estimates have been boosted by using ML.
Table 3: Forecast on the future precision of optical GRB cosmology.
@@
 [^7] Number of optical GRBs with a plateau and redshift needed to reach the precision of $\Omega_{M}$ as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia). This is an extension of the table provided in Dainotti et al. (2022). The term n=1 indicates that the error bars on the GRB plateau parameters are full width, while n=2 indicates that the error bars are halved. The ”LCR” indicates that the estimates for the number of GRBs have been boosted by using LCR, while ”ML” indicates that the estimates have been boosted by using ML.
@@ 5.6 Applicability to future missions
It is important to note that all the methodologies spelled out here can be applied not only to GRBs observed by Swift, but also to any GRB for which we can observe the plateau emission. Thus, we envision this method can be applied to GRBs observed by Space-based multi-band astronomical Variable Objects Monitor (SVOM,Wei et al. (2016)), the Transient High-Energy Sky and Early Universe Surveyor (THESEUS) mission (Amati et al., 2018), and the Einstein Probe mission(Yuan et al., 2015). Surely, with the new data from other satellites, one needs to be cautious about how the variables’ parameter spaces will change compared to the space of the Swift variables and how much this will impact the prediction. Properly leveraging the power of these new satellites will require a re-training of the ML models. However, the base methodology described here and in our previous works (Dainotti et al., 2024c, b) should still produce reliable results.
@@ 5.7 Redshift inference with the web app
Our web app has various functionalities, but its main focus is inferring the redshift of GRBs without $z_{obs}$. Users have the flexibility to upload a file containing the predictors discussed in this work (see Sec. 4.2.1) or manually enter the parameters of a single GRB for redshift estimation. This interface is presented in Fig. 18. This module of the web app also allows the user to obtain MICE imputations if their data contains missing values. To do so, they select the ‘Apply MICE’ button. After providing the data, the web app initiates the back-end ML codes, the results and performance of which have been presented in this work. Upon successful execution, the user is presented with individual redshift values for each GRB in a tabular manner, which can also be downloaded as a text file. This table displays the GRB IDs along with $z_{pred}$ and its associated uncertainties. Furthermore, if a user already has their custom model, trained from modules such as Model Comparision and SuperLearner, then they can use that model for predictions. For this, they need to select the ”Do you have modeled data?” option. However, the users can use the model trained on our data without necessarily providing their own training set. Finally, comprehensive documentation is available on the web app in the ‘Documentation’ module to help users navigate.
Figure 18: Redshift Estimator, the homepage of the app, depicted with the options of choosing MICE, uploading a file, entering parameters manually, and using custom models.
@@ 6 Discussion
The findings from Dainotti et al. (2024b) demonstrate the promise of using machine learning for redshift estimation, particularly in overcoming the limited number of GRBs with precise redshift measurements. In this work, we have expanded both the methodology and the sample sizes for the training and generalization data sets. Providing redshift predictions for the generalization set will help future population studies address the GRBFR and SFR discrepancy and other relevant issues. In the future, the $z_{pred}$ values obtained here can also be utilized to explore the characteristics of the central engine in GRB events, focusing on GRBFR, LF, and the structure of LCs. We can also use GRB LCs obtained from reconstruction, which some of us did in Dainotti et al. (2023b). We demonstrated the effectiveness of this reconstruction technique in reducing the errors on plateau parameters by an average of 30%-40%.
Another point we wish to highlight here is that even though we have the largest sample of LGRBs with plateau and redshift, this is still at least an order of magnitude less than the data required to reliably train more complex ML models, such as neural networks. Indeed, without including the plateau, Swift XRT observed 1420 GRBs up to December 2024, which is too small to train complex ML models properly. In some of our previous works, using Active Galactic Nuclei data (Dainotti et al., 2021a; Narendra et al., 2022; Gibson et al., 2022), we have shown that SuperLearner picks different, more complex models (e.g., XGBoost, EARTH) for the larger sample size. Similarly, we tested multiple ML models to find the best set of models that can give accurate redshift predictions with the current data set. The three chosen models performed the best during our analysis. Given the size of our data set, parametric and semi-parametric functions such as GLM and GAM perform better. This fact is further highlighted in the Appendix in Fig. 20, where the weight assigned by SuperLearner to the model EARTH is lower than the weights assigned to the GAM, GLM, and RF models. This is in contrast with the results presented in Narendra et al. (2022) where Earth was one of the best models. Furthermore, to test whether the performance of these models is maintained as the data set increases, we have also performed a MC simulation (see Sect. 5.4). From those simulations we can say that our models will be reliable for up to three years, after which we will require them to be retrained with different models and training data. Thus, based on our analysis, these chosen models are best suited for the data at hand.
6.1 LGRBs versus SGRBs